OpenStudy (anonymous):

you invest $6000 in two accounts paying 7% and 8% annual interest, respectively. If the total interest earned for the year was $440, how much was invested at each rate?

OpenStudy (anonymous):

please help!

OpenStudy (mertsj):

Let x = the amount invested at 7% Then 6000-x was invested at 8% The interest on the x dollars is .07x the interest on the 6000-x dollars is .08(6000-x) The total interest is 440 so we have: \[.07x+.08(6000-x)=440\]

OpenStudy (mertsj):

We should probably multiply by 100 to get rid of all the decimals.

OpenStudy (mertsj):

\[7x+8(6000-x)=44000\]

OpenStudy (mertsj):

Solve that equation.

OpenStudy (anonymous):

7x+48000-8x=44000

OpenStudy (anonymous):

?

OpenStudy (mertsj):

Combine the like terms.

OpenStudy (anonymous):

-1x+48000=44000

OpenStudy (mertsj):

subtract 48000 from both sides.

OpenStudy (anonymous):

i got 4000 as my answer but dont knw what that is for

OpenStudy (mertsj):

What did we say x stands for in the beginning when we defined the variables/

OpenStudy (anonymous):

ohhh duh

OpenStudy (anonymous):

thank you:D

OpenStudy (anonymous):

i have more problems can u help?

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

Posssibly if they are not too hard.

OpenStudy (anonymous):

all math is hard for me! haha

OpenStudy (anonymous):

the sum of two numbers is 13. If one number is subtracted from the other the result is 3 find the nmbers.

OpenStudy (mertsj):

x+y=13 x-y=3

OpenStudy (mertsj):

Can you solve that system?

OpenStudy (anonymous):

they cancel dont they?

OpenStudy (mertsj):

Yes. If you add the 2 equations, the y's cancel and you have 2x=16

OpenStudy (anonymous):

8